# LinkAlign: Scalable Schema Linking for Real-World Large-Scale Multi-Database Text-to-SQL

### v2

@@ 5.1 Experiemental Setup
Dataset. We evaluate our method performance of schema linking on the SPIDER, BIRD and AmbiDB datasets, and the ability to adapt existing Text-to-SQL models to real-world environments on the SPIDER 2.0 benchmark. We provide more details about SPIDER, BIRD and SPIDER 2.0 in Appendix  E, and the construction of the AmbiDB dataset in Appendix  D.
Baselines.We compare our method with the following baselines.
$\bullet$ DIN-SQL Pourreza and Rafiei (2024) introduces a prompt-driven schema linking method with a single LLM call which integrates a chain-of-thought strategy to enhance reasoning.
$\bullet$ PET-SQL Li et al. (2024b) leverages LLMs’ SQL generation capabilities to generate preliminary SQL (PreSQL) statements for deriving schema references.
$\bullet$ MAC-SQL Wang et al. (2024) employs a Selector agent to identify the minimal relevant schema subset with a single LLM call.
$\bullet$ MCS-SQL Lee et al. (2024) employs a two-step process for table and column linking, utilizing multiple prompts and random shuffling to enhance recall and robustness.
$\bullet$ RSL-SQL Cao et al. (2024) introduces a bidirectional schema linking approach that integrates both forward and backward linking to streamline the database schema.
@@
Hyperparameter. The open-sourced text-embedding model bge-large-en-v1.5 is applied to convert database schema metadata and queries into vectors. We set the top-K of retrieval size at 5 and we use GLM-4-air model for experiments.
@@ 5.2 Main Results
Results on Multi-Databases. As shown in Table  1, our method achieves state-of-the-art Locate Accuracy (LA) on the SPIDER, BIRD, and AmbiDB datasets, with scores of 86.4%, 83.4%, and 69.4%, respectively, demonstrating the framework’s effectiveness in mapping data requirements to the target database. Unlike baseline methods, our framework introduces the "Reduce Noise by Response Filtering" step that filters out irrelevant database schemas, reducing error 2 and maintaining high LA accuracy. This step also benefits the model by eliminating the need to choose between similar candidate schemas, allowing it to identify the required tables and columns more accurately. Experimental results confirm this: our framework achieves state-of-the-art Exact Match (EM) performance across all three datasets, surpassing the baseline models by 23.6%, 1.8%, and 37.3%, respectively. The presence of error 1,2 in multi-database contexts makes it difficult for the model to avoid blending unrelated schemas, leading to challenges in accurately recalling relevant tables and columns. Especially as the database size increased, we observed a significant decrease in the recall rate of all methods on the AmbiDB dataset, further demonstrating that our proposed dataset exacerbated the challenge.
@@
Results on Single Database. We further compare and evaluate the ability of different methods to identify correct tables and columns in a large-scale database. As shown in Table  2, our Agent method achieves state-of-the-art performance across all datasets in terms of the EM evaluation metric. After eliminating interference from irrelevant database schemas, the recall rate for all methods significantly improved compared to the results in Table  1. Compared to the baseline models, the Agent method achieves the highest recall rates on both the Spider-dev and Bird-dev datasets. This demonstrates that our method exhibits superior performance and robustness when the inference capabilities of large models are limited. Although, on the AmbiDB dataset, our method’s recall rate lags behind MCS-SQL (88.5%) and RSL-SQL (88.3%), our method outperforms these models by 21.3% and 7.4%, respectively, in Precision. This indicates that our method maintains high recall while minimizing irrelevant noise. Overall, considering all metrics, our method demonstrates excellent performance and robust schema-linking capabilities.
@@
Results on Advanced Spider 2.0-lite Benchmark. To convincingly demonstrate the effectiveness of our method, we conducted tests on Spider 2.0 Lei et al. (2024), an open-source benchmark which is tailored for evaluation on real-world challenges. Our submission results are shown in the Figure 1. Notably, our method using Qwen-PLUS significantly outperforms other models like DailSQL, CHESS, and DIN-SQL using GPT-4o, highlighting the effectiveness of our approach. Furthermore, our approach, LinkAlign+DeepSeek-v3 ranks overall 4th place across all sequences without introducing test-time scaling, in contrast to the top three methods which all employ OpenAI O1 and O3 models. This underscores the potential of our method for further improvement in the future.
@@ 5.3 Ablation Study
We conducted a ablation study to investigate incremental impact of two core step of LinkAlign framework.The results are presented in Tables  3 , demonstrated that each step of proposed methods contributes to the SOTA performance on benchmark datasets.
Impact of Query Rewriting. User queries often fail to align with their required schema, leading to inefficient retrieval. As shown in Figure  4, introducing Query Rewriting step reduced error 1 rates of two strategies by 6.9% and 10.8%, respectively. This demonstrates this step effectively eliminates ambiguity and improves retrieval recall of target database.Compared to Pipeline , the Agent strategy shows more significant benefits, suggesting LLMs reflextion more effectively aligns the query with the schema. However, this step also introduces more irrelevant database schemas, which increases the challenge of database-locate and leads to a rise in error 2 by 5.7% and 8.4%, respectively.Since the slope of the negative impact from error 2 is lower than the positive gain from solving error 1, the Locate Accuracy metric shows an overall improvement. Compared to the Spider benchmark, the LA improvement on the AmbiDB dataset is more pronounced, indicating this step plays a more critical role when ambiguity in queries increases. For simple query scenarios, it may be necessary to balance the positive gains and negative impacts of query rewriting to maximize the LA benefits.
Impact of Response Filtering. As shown in Figure  4, although Query Rewriting introduces many irrelevant databases, the Response Filtering step reduces the error 2 rate of the Agent strategy by 10.8%, effectively mitigating the negative impact of the previous step. Compared to the Pipeline strategy, the Agent strategy benefits more significantly, indicating that this step plays a more critical role within the Agent strategy. Filtering allows LLMs to focus on single-target database schema more narrowly, making it easier to accurately identify the most relevant schema. As shown in Table  3, Response Filtering effectively improves both the EM and Recall metrics for both strategies. Although selecting the wrong database directly causes failure in Schema Linking task, Response Filtering significantly improves schema linking performance by mitigating the negative impact of error 3,4.
@@ 7 Limitations
Appendix A Error Analysis
@@
Appendix B Single-Prompt Pipeline Strategy
@@
Appendix C Effectiveness in Single-Datbase Setting
@@
Appendix D AmbiDB Dataset Construction
@@ D.2 Data Filter
Appendix E Supplementary Experimental Setup
@@
Appendix F SPIDER 2.0-lite Evaluation
We evaluate our method on the SPIDER 2.0-lite Benchmark. In the experiments, we applied LinkAlign to the DIN-SQL base model to adapt it to real-world large-scale query scenarios. Our method rank highest excluding those using long CoT reasoning LLMs. The results demonstrate that our method has a significant impact.
Figure 6: Performance of LinkAlign framework on SPIDER2.0-lite benchmark.

### v3

@@ 4.1 Experiemental Setup
Dataset. We evaluate our method performance of schema linking on the SPIDER, BIRD and AmbiDB datasets, and the ability to adapt existing Text-to-SQL models to real-world environments on the SPIDER 2.0 benchmark. We provide more details about SPIDER, BIRD and SPIDER 2.0 in Appendix  F, and the construction of the AmbiDB dataset in Appendix  E.
Baselines. We compare our method against multiple LLM-based schema linking methods. DIN-SQL Pourreza and Rafiei (2024) employs a prompt-driven approach with a single LLM call and a chain-of-thought strategy to improve reasoning. PET-SQL Li et al. (2024b) generates preliminary SQL to infer schema references. MAC-SQL Wang et al. (2024) uses a Selector agent to identify minimal relevant schema subsets. MCS-SQL Lee et al. (2024) applies a two-step table and column linking process with multiple prompts and random shuffling for robustness. RSL-SQL Cao et al. (2024) adopts a bidirectional linking strategy that combines forward and backward schema linking.
@@
Implementations. The open-sourced text-embedding model bge-large-en-v1.5 is applied to convert database schema metadata and queries into vectors. We set the top_k of the retrieval size at 5 and adaptively adjust turn_n according to the database size. We use GLM-4-air model for schema linking and DeepSeek-V3, R1 and Qwen-72B for "end-to-end" Text-to-SQL evaluation. We further developed a versatile Text-to-SQL development and evaluation tool, enabling multi-task concurrent calls via configuration files, thereby supporting subsequent experimental testing.
@@ 4.2.1 Schema Linking Performance
Multi-Databases Results. As shown in Table  2, our method achieves the highest Locate Accuracy (LA) on the SPIDER, BIRD, and AmbiDB datasets, with scores of 86.4%, 83.4%, and 69.4%, respectively, demonstrating the effectiveness in mapping data requirements to the target database. A key contributor to this improvement is the introduction of the Response Filtering step, which mitigates Error 2 by eliminating irrelevant database schemas. Additionally, our framework achieves the highest Exact Match (EM) performance across all three datasets, outperforming baseline models by margins of 23.6%, 1.8%, and 37.3%, respectively.The presence of error 1,2 in multi-database contexts makes it difficult to avoid blending unrelated schemas, leading to challenges in accurately recalling relevant tables and columns. Especially as the database size increased, we observed a significant decrease in the recall rate of all methods on the AmbiDB dataset, further demonstrating that our proposed dataset exacerbated the challenge.
@@
Single-Database results. We further compare and evaluate the ability of different methods to identify correct tables and columns in a large-scale database. As shown in Table  3, our Agent method achieves state-of-the-art performance across all datasets in terms of the EM evaluation metric. After eliminating interference from irrelevant database schemas, the recall rate for all methods significantly improved compared to the results in Table  2. Compared to the baseline models, the Agent method achieves the highest recall rates on both the Spider-dev and Bird-dev datasets. This demonstrates that our method exhibits superior performance and robustness when the inference capabilities of large models are limited. Although, on the AmbiDB dataset, our method’s recall rate lags behind MCS-SQL (88.5%) and RSL-SQL (88.3%), our method outperforms these models by 21.3% and 7.4%, respectively, in Precision. This indicates that our method maintains high recall while minimizing irrelevant noise. Overall, considering all metrics, our method demonstrates excellent performance and robust schema-linking capabilities.
4.2.2 Text-to-SQL Performance
Spider 2.0-lite Results. To convincingly demonstrate the effectiveness of the framework, we conducted tests on the Spider 2.0-Lite benchmark Lei et al. (2024), which simulates real-world challenges by significantly increasing the number of schemas. As shown in Table 1, we achieved the new SOTA score of 33.09% applying LinkAlign to the DIN-SQL method of lowest rank. In particular, our method achieves performance comparable to the existing baseline like ReFoRCE and Spider-Agent using purely open-source LLMs. The results highlight the framework’s effectiveness in enhancing Text-to-SQL performance by improving schema linking in large-scale database environments.
Small-scale database results. We assess the framework ability to generalize improved schema linking to smaller-scale databases by evaluating it on the Spider and Bird dev set. To further assess generalization across diverse LLMs, we employed two open-source models, Deepseek and Qwen, with significantly different parameter sizes of 671B and 72B, respectively. Given that the limited database schemas would not exceed the model’s context and minor redundant schema noise would not impact the LLMs’ attention significantly, we only utilized a simplified LinkAlign framework by excluding Step One and Step Two. The results show that Execution Accuracy gains of 6.7% on Spider and 7.2% on Bird, demonstrating that improved schema linking enhances SQL generation significantly.
Approach EX (%)
DIN-SQL + GPT-4 82.8
MAC-SQL + GPT-4 86.8
DAIL-SQL + GPT-4 86.6
MCS-SQL + GPT-4 89.5
LinkAlign+DeepSeek-V3(671B) 88.9
LinkAlign + Qwen(72B) 86.8
Table 4: Comparison of different methods on Spider-dev dataset. indicates method using a simplified LinkAlign framework without Step One and Step Two.
Approach EX (%)
DIN-SQL + GPT-4 50.7
MAC-SQL + GPT-4 59.4
DAIL-SQL + GPT-4 54.8
MCS-SQL + GPT-4 63.4
RSL-SQL + GPT-4 67.2
LinkAlign+DeepSeek-V3(671B) 57.5
LinkAlign + Qwen(72B) 53.4
Table 5: Comparison of different methods on Bird-dev dataset. indicates method using a simplified LinkAlign framework without Step One and Step Two.
4.3 Runtime Efficiency
We assessed the average runtime of each step in LinkAlign using samples from the Spider 2.0-lite dataset. The results show that pipeline mode is more efficient and better suited for latency-sensitive scenarios, while agent mode offers improved performance when accuracy is prioritized. This flexibility enables users to adapt LinkAlign to different application needs.
Approaches S1 Time (s) S2 Time (s) S3 Time (s)
Pipeline 9.02 2.94 1.67
Agent 30.90 26.23 13.46
Table 6: Average time for each step of the framework.
@@ 4.4 Ablation Study
We conducted an ablation study to examine the incremental impact of the two core steps in the LinkAlign framework. We exclude Step 3 from consideration, as schema parsing is often considered fundamental to schema linking and must be retained. As shown in Table 7, each step contributes to achieving SOTA performance on the benchmark.
Impact of Query Rewriting. User queries often misaligned with the target schema, leading to retrieval inefficiency. As shown in Figure 3, adding Query Rewriting reduces Error 1 by 6.9% and 10.8% for two mode, improving recall by resolving ambiguity. The Agent mode benefits more than Pipeline, indicating LLM reflection better aligns queries to schemas. However, this step also introduces irrelevant schemas, increasing Error 2 by 5.7% and 8.4%, which complicates database localization. Despite this, the net effect is positive: Locate Accuracy improves overall. The improvement is more notable on the AmbiDB dataset than on Spider, showing query rewriting is especially important with higher ambiguity. For simple queries, balancing gains and drawbacks of rewriting is recommended to optimize Locate Accuracy.
Impact of Response Filtering. Figure 3 shows that although Query Rewriting introduces irrelevant databases, Response Filtering reduces Error 2 in the Agent mode by 10.8%, effectively offsetting this negative effect. The Agent mode gains more than Pipeline, highlighting the filtering step’s critical role in narrowing LLM focus to the correct schema. As Table 7 demonstrates, Response Filtering improves both EM and Recall for both strategies. Despite wrong database selection causing Schema Linking failures, filtering significantly boosts schema linking by mitigating Error 3 and 4.
@@ 6 Limitations
Appendix A Related Work
Appendix B Error Analysis
@@
Appendix C Single-Prompt Pipeline Strategy
@@
Appendix D Effectiveness in Single-Datbase Setting
@@
Appendix E AmbiDB Dataset Construction
@@ E.2 Data Filter
Appendix F Supplementary Experimental Setup